A sack slides off the ramp, shown in Fig. 3, with a horizontal velocity of 12 m/s. If the height of the ramp is 6 m from the floor, determine the time needed for the sack to strike the floor and the range R where sacks begin to pile up.
1
Expert's answer
2020-08-24T11:57:26-0400
"t=\sqrt{\frac{2h}{g}}=\sqrt{\frac{2(6)}{9.8}}=1.1\ s"

"R=vt=v\sqrt{\frac{2h}{g}}=12\sqrt{\frac{2(6)}{9.8}}=13.3\frac{m}{s}"